OpenStudy (anonymous):

Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal.

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8

OpenStudy (anonymous):

@Hero @ParthKohli

hero (hero):

@memecheaa, is that the full instructions for the question? I assume this is related to I = PRT problems.

OpenStudy (anonymous):

Yes. it is You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% Using the function A(t)=P(1+rn)nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down.

hero (hero):

Okay, there's the rest of it. Is the function \(A(t) = P(1 + rn)^{nt}\) ? How far have you gotten with solving this?

OpenStudy (anonymous):

@Hero: I have gotten: A(t)=39143(1+0.03/12)^12t

hero (hero):

Okay, yes, that makes more sense. The formula should be \(A(t) = P\left(1 + \dfrac{r}{n}\right)^{nt}\)

OpenStudy (anonymous):

Soooo 39143t=10,000(1=0.03/12)^(12t)

OpenStudy (anonymous):

?*

hero (hero):

Actually, I re-read it. It clearly says the principle should be the price of the vehicle.

OpenStudy (anonymous):

Was i right the first time?

hero (hero):

So what you had before is correct. \(10,000 = 39143\left(1 + \dfrac{0.03}{12}\right)^{12t}\)

hero (hero):

That probably explains why they said t will be negative.

OpenStudy (anonymous):

It's not 10,000t?

hero (hero):

It's definitely not 10,000t.

OpenStudy (anonymous):

Oh.. Okay . \[1+.0025=1.0025\] then i multiply it to 39143 right?

hero (hero):

Basically, all you have to do is just solve for t.

hero (hero):

But you can't multiply 39143 with 1.0025 because the 1.0025 is raised to an exponent.

hero (hero):

The first step is to divide both sides by 39143

hero (hero):

\(\dfrac{10,000}{39143} = \left(1.0025\right)^{12t}\)

OpenStudy (anonymous):

Ohhhhhhhh .. now do we raise 1.0025?

OpenStudy (anonymous):

1.0025^(12)(1)= 1.0304

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Try to take log of both sides

hero (hero):

The next step is to take logs of both sides like what @geerk42 said.

hero (hero):

It's probably a good idea to reduce \(\dfrac{10,000}{39143}\) before doing so.

OpenStudy (anonymous):

@Hero thank you soooo much! :) . can you help me with one more please?

hero (hero):

Post it as a new question.